A stage has a width that is 6 feet less than it’s length. The area of the stage is 720 ft2. What is the length of the stage?
Found 2 solutions by Alan3354, Cromlix:
Answer by Alan3354(69443)

A stage has a width that is 6 feet less than it’s [sic] length. The area of the stage is 720 ft2. What is the length of the stage?
----------------
Find a pair of factor of 720 that differ by 6.
PS it's = it is

Answer by Cromlix(4381)

Hi there,
Make the length = x
Width = x - 6
Area of rectangle
= length * width (* = times)
720 = x(x - 6)
Multiply out
720 = x^2 - 6x
Rearrange into ax^2 + bx + c = 0 form
x^2 - 6x - 720 = 0
Factorize
(x + 24)(x - 30) = 0
x + 24 = 0
x = -24 (disregard as -ve)
x - 30 = 0
x = 30
Length = 30 feet
Width = 24 feet.
Hope this helps :-)